An airplane flies from City A in a straight line to City B, which is 60 kilometers north and 150 kilometers west of City A. How far does the plane fly? (Round your answer to the nearest kilometer.)

Answer by geetha_rama(94)

City B is 60 Km North and 150Km west It is a right angled triangle where c is the hypotenuse which is the diastance travelled by the plane
c^2 = (sqrt(60^2+150^2))
= (sqrt(3600+22500))
= (sqrt(26100))
= 161.55
Aeroplane would have travelled 162 Km